# ASTREA: Introducing Agentic Intelligence for Orbital Thermal Autonomy

### v1

@@ Abstract
This paper presents ASTREA, the first agentic system deployed on flight-heritage hardware (TRL 9) for autonomous spacecraft operations. Using thermal control as a representative use case, we integrate a resource-constrained Large Language Model (LLM) agent with a reinforcement learning controller in an asynchronous architecture tailored for space-qualified platforms. Ground experiments show that LLM-guided supervision improves thermal stability and reduces violations, confirming the feasibility of combining semantic reasoning with adaptive control under hardware constraints. However, on-orbit validation aboard the International Space Station (ISS) reveals performance degradation caused by inference latency mismatched with the rapid thermal cycles characteristic of Low Earth Orbit (LEO) satellites. These results highlight both the opportunities and current limitations of agentic LLM-based systems in real flight environments, providing practical design guidelines for future space autonomy.
@@ 1 INTRODUCTION
Recent advances in robotics and artificial intelligence have accelerated the possibilities for autonomy, particularly in space exploration missions. However, AI could potentially transform the way lunar missions and Earth-orbiting missions are operated and monitored, especially in scenarios requiring real-time decision-making as an integral part of mission operations. This technological evolution represents a paradigmatic shift toward space systems capable of operating with minimal human intervention, particularly critical in environments where communication delays preclude direct ground supervision. Numerous autonomous control initiatives have been documented, ranging from assistance in traditional subsystems to autonomous capture and landing mission planning. These developments have demonstrated the technical feasibility of delegating critical decisions to automated systems, establishing the foundation for the next generation of fully autonomous space missions. Recently, the emergence of Large Language Models (LLMs) and their integration into the new agentic paradigm have added a new dimension to autonomous decision-making capabilities, enabling the interpretation of nuances, ambiguous contexts, and non-deterministic decision-making based on objectives or historical data, approaching human-like behavior in decision processes. This semantic reasoning capability opens previously unimaginable possibilities for space autonomy, where contextual interpretation and adaptability are fundamental to mission success. However, these large models also require substantial hardware resources in their most powerful versions, which conflicts with embedded environments that must maintain radiation tolerance and operate under severe constraints regarding power consumption, size, and temperature. This dichotomy between desired capabilities and the physical limitations of space hardware represents one of the primary challenges for the practical implementation of agentic systems in flight applications. Thermal control constitutes a critical subsystem that perfectly exemplifies these challenges, as it must maintain the operational integrity of all electronic components while managing limited computational resources in real-time. The inherent complexity of balancing variable thermal loads with dynamic processing demands requires a level of decision-making sophistication that has traditionally relied on pre-programmed rules and ground-based supervision. This work addresses this gap by implementing an agentic system on flight-qualified hardware (TRL 9) currently operating in the IMAGIN-e[1] mission, where an LLM agent provides contextual recommendations for thermal control parameters while maintaining operational independence for safety-critical applications. Beyond ground validation, the system was also deployed in on-orbit experiments aboard the International Space Station (ISS), representing, to the best of our knowledge, the first demonstration of an agentic LLM-based supervisory system executed in a real flight environment.
@@ 3.2 LLM Agent
Initial preliminary experiments revealed LLM response times of approximately 10 minutes in certain instances. Considering these durations, alongside the time required for the RL model to adapt to new alpha values and cool-down periods between episodes, a 60-minute time window was established for collecting episode summaries before triggering alpha value recommendations.
This window size ensures sufficient data collection for meaningful LLM recommendations while maintaining reasonable system responsiveness for thermal control applications.
@@ 3.5.1 Ground Laboratory Setup
The ground experiments were conducted in a mini-rack configuration equipped with active cooling via fans. Ambient temperature was regulated for 10 hours during the day but remained unconditioned overnight, introducing natural thermal fluctuations. The SoC operated under continuous computational stress, with 15 cores fully loaded and core 0 reserved for agentic orchestration.
@@ 3.6.1 Baseline: SAC with adaptive $\alpha$
The baseline system employs the same RL-agent architecture described in Section 3.3, but uses the default adaptive $\alpha$ scheduling mechanism from Stable Baselines3 [14]. This implementation automatically adjusts the entropy coefficient during training according to the internal heuristics of the algorithm, without external intervention or episode-level performance feedback 5.
@@ 4.2 On-Orbit Experiments (ISS)
ASTREA was subsequently deployed aboard the International Space Station (ISS) to validate feasibility in a true operational environment (See 3.5.2). In this experiment, the agentic system exhibited shorter episode durations than the RL-only baseline, but also achieved a reduction in the number of thermal violations. CPU usage showed a slight increase when the LLM-agent was active (Table 2).
Table 2: Representative results from on-orbit experiments aboard the ISS.
@@
Avg. Episode Duration $\mathbf{224.26\pm 163.81}$ $146,60\pm 71.04$
Thermal Violations $204\pm 201.72$ $\mathbf{184.33\pm 69.47}$
Avg. CPU Usage (%) $18.84\pm 13.01$ $\mathbf{19.72\pm 17.43}$
The mixed performance profile can be partially explained by a temporal mismatch between the LLM decision cycle and the orbital thermal dynamics, combined with the stabilizing effect of the agentic policy on thermal control. Unlike terrestrial experiments, where ambient temperature evolves gradually and predictably, the ISS undergoes rapid and extreme thermal transitions due to its 90-minute orbital period. As the station circles the Earth, it experiences approximately 45 minutes of direct sunlight followed by 45 minutes of eclipse, resulting in sharp temperature swings across its external surfaces.
In this experiment, the analysis window provided to the LLM was shortened to 15 minutes to approximate this orbital cadence. However, as illustrated in Figure 7, the LLM required up to 8 minutes to generate a recommendation in the current configuration. During such an interval, the ISS may already undergo a complete sun/eclipse transition, rendering the suggested $\alpha$ value outdated at the moment of application. Consequently, some updates introduced exploratory behavior precisely during phases that demanded conservative exploitation, while in other cases the agentic adjustments reduced the magnitude and frequency of thermal violations.
Despite the mixed outcomes observed in orbit, the successful deployment and execution of the agentic system aboard the ISS validates the feasibility of the onboard architecture under real flight conditions. This operational confirmation represents a critical milestone for future agentic system implementations in space environments.
@@ 4.3 Comparative Discussion
The discrepancy between ground and on-orbit results underscores an important limitation of agentic LLM-based systems in flight conditions: when the environment’s dominant timescale (e.g., orbital sun/eclipse transitions) is comparable to or shorter than the LLM inference latency, semantic supervision can become counterproductive. However, the observed reduction in thermal violations suggests that, even under latency constraints, agentic policies can contribute to improved stability in certain operational metrics. This finding highlights a general principle for the deployment of LLMs in safety-critical space systems: they are unsuitable for direct real-time intervention unless the latency of decision-making is significantly smaller than the dynamics of the controlled process.
Nevertheless, these mixed results are highly informative. They provide empirical evidence that complements the successful ground experiments and establish practical design guidelines.
@@ 5 Implications for Future Space Autonomy
The experimentation confirms that the integration of linguistic models for semantic supervision of other processes is feasible in flight-qualified equipment within the transformative New Space era. Agentic systems are thus positioned as a viable and, given their inherent characteristics, transformative approach to integrating control logic into traditional subsystems. The deployment of ASTREA demonstrates how agentic systems can serve as supervisory layers for traditional control architectures. By combining linguistic reasoning with adaptive control, ASTREA exemplifies a scalable approach to onboard autonomy, capable of supporting fault detection, mission planning, and subsystem coordination.
@@ 6 Conclusions and Future Work
This study addresses the utilization of agentic systems in real flight equipment to provide an advanced layer of pseudo-reasoning at the edge. The research evaluates the utility of small linguistic models, which despite their contextual understanding limitations, remain feasible given current hardware constraints in flight equipment, deployed at the edge for autonomous control applications.
ASTREA represents a pioneering implementation of agentic autonomy in flight-qualified hardware. By integrating a quantized LLM with a reinforcement learning agent, ASTREA establishes a framework for semantic supervision and adaptive control in space systems. Its modular design and asynchronous architecture offer a scalable blueprint for future multi-agent deployments.
To validate this approach, we experimented with the integration of an LLM-agent with a reinforcement learning agent previously employed in earlier studies for autonomous thermal control of an edge processing node through the management of available processing resources. While the models present inherent limitations, our findings demonstrate that through the application of various optimization techniques, their utilization becomes viable, yielding clear improvements in the thermal control agent’s performance, during the cold start phase, by modifying its behavior based on obtained results. This highlights the potential of linguistic models for long-term context analysis, efficiently complementing the capabilities of reinforcement learning models in decision-making processes.
The present study employed a single LLM agent with significant response time limitations. This constraint would become even more critical in multi-agent systems where bottlenecks could easily emerge, making it worthwhile to explore these limitations in future research.
Since the flight equipment utilized lacks an accelerator for optimal LLM model execution, deployment occurred on the CPU alongside other concurrent processes. While functionality has been demonstrated, the incorporation of a space-qualified accelerator would fundamentally transform system performance, potentially enabling the use of more complex models or alternative execution designs such as shorter analysis windows and more frequent evaluations.
In the analyzed use case, the LLM agent suggests entropy coefficient values for a reinforcement learning model. However, numerous other parameters within such models could also be analyzed by the LLM agent, including the possibility of adjusting the reward function to shape the behavioral objectives of the RL agent, as explored by Navarro T. et al. [6].
Moreover, the capabilities of LLMs are not limited to fine-tuning existing parameters; their potential extends to higher-level control mechanisms. Thanks to their ability to interpret broader contexts, LLM agents could monitor instrument degradation throughout a mission and adapt usage strategies accordingly. They might also detect periodic patterns in environmental conditions based on orbital position and dynamically select among reinforcement learning models with varying characteristics—such as more conservative or more aggressive behaviors. Furthermore, these models could be combined or switched in real time depending on evolving operational priorities. For instance, the system might prioritize energy efficiency over computational performance during specific mission phases, or vice versa, depending on the mission’s current objectives.
Finally, this study selected a general-purpose model that may lack sufficient thermal control domain knowledge. Repeating the experiment with a model specialized in thermal control would be highly interesting, as enhanced task-specific knowledge could enable more grounded analysis without requiring extensive prompt engineering to condition its pseudo-reasoning capabilities.
In summary, ASTREA proved effective in the lab and, in orbit, reduced thermal violations and increased CPU utilization in line with mission goals, although episode durations were shorter. These results confirm the feasibility of LLM‑based supervision in flight and provide concrete design criteria for future spaceborne autonomous systems.

### v2

@@ Abstract
This paper presents ASTREA, the first agentic system executed on flight-heritage hardware (TRL 9) for autonomous spacecraft operations, with on-orbit operation aboard the International Space Station (ISS). Using thermal control as a representative use case, we integrate a resource-constrained Large Language Model (LLM) agent with a reinforcement learning controller in an asynchronous architecture tailored for space-qualified platforms. Ground experiments show that LLM-guided supervision improves thermal stability and reduces violations, confirming the feasibility of combining semantic reasoning with adaptive control under hardware constraints. On-orbit validation aboard the ISS initially faced challenges due to inference latency misaligned with the rapid thermal cycles of Low Earth Orbit (LEO) satellites. Synchronization with the orbit length successfully surpassed the baseline with reduced violations, extended episode durations, and improved CPU utilization. These findings demonstrate the potential for scalable agentic supervision architectures in future autonomous spacecraft.
@@ 1 INTRODUCTION
Recent advances in robotics and artificial intelligence have accelerated the possibilities for autonomy, particularly in space exploration missions. However, AI could potentially transform the way lunar missions and Earth-orbiting missions are operated and monitored, especially in scenarios requiring real-time decision-making as an integral part of mission operations. This technological evolution represents a paradigmatic shift toward space systems capable of operating with minimal human intervention, particularly critical in environments where communication delays preclude direct ground supervision.
Numerous autonomous control initiatives have been documented, ranging from assistance in traditional subsystems to autonomous capture and landing mission planning. These developments have demonstrated the technical feasibility of delegating critical decisions to automated systems, establishing the foundation for the next generation of fully autonomous space missions.
Recently, the emergence of Large Language Models (LLMs) and their integration into the new agentic paradigm have added a new dimension to autonomous decision-making capabilities, enabling the interpretation of nuances, ambiguous contexts, and non-deterministic decision-making based on objectives or historical data, approaching human-like behavior in decision processes. This semantic reasoning capability opens previously unimaginable possibilities for space autonomy, where contextual interpretation and adaptability are fundamental to mission success.
However, these large models also require substantial hardware resources in their most powerful versions, which conflicts with embedded environments that must maintain radiation tolerance and operate under severe constraints regarding power consumption, size, and temperature. This dichotomy between desired capabilities and the physical limitations of space hardware represents one of the primary challenges for the practical implementation of agentic systems in flight applications.
Thermal control constitutes a critical subsystem that perfectly exemplifies these challenges, as it must maintain the operational integrity of all electronic components while managing limited computational resources in real-time. The inherent complexity of balancing variable thermal loads with dynamic processing demands requires a level of decision-making sophistication that has traditionally relied on pre-programmed rules and ground-based supervision. This work addresses this gap by implementing an agentic system on flight-qualified hardware (TRL 9) currently operating in the IMAGIN-e[1] mission, where an LLM agent provides contextual recommendations for thermal control parameters while maintaining operational independence for safety-critical applications. Beyond ground validation, the system was also deployed in on-orbit experiments aboard the International Space Station (ISS), representing, to the best of our knowledge, the first demonstration of an agentic LLM-based supervisory system executed in a real flight environment.
@@
The paper proceeds as follows: Section 2 reviews the state of the art in LLM-based intelligence autonomy for space exploration. Section 3 details the experimental design and implementation of ASTREA, including the LLM and RL agents, hardware platform, and evaluation environments. Section 4 presents the results from ground and on-orbit experiments, along with a comparative analysis against the baseline. Section 5 discusses the implications for future space autonomy, and Section 6 concludes with key findings and directions for future work.
@@ 3.2 LLM Agent
Initial preliminary experiments revealed LLM response times of approximately 10 minutes in certain instances. Considering these durations, alongside the time required for the RL model to adapt to new alpha values and cool-down periods between episodes, a time window was established for collecting episode summaries before triggering alpha value recommendations.
The specific window size was selected based on the operational environment to ensure sufficient data collection for meaningful LLM recommendations while maintaining reasonable system responsiveness for thermal control applications.
@@ 3.5.1 Ground Laboratory Setup
The ground experiments were conducted in a mini-rack configuration equipped with active cooling via fans. Ambient temperature was regulated for 10 hours during the day but remained unconditioned overnight, introducing natural thermal fluctuations. The SoC operated under continuous computational stress, with 15 cores fully loaded and core 0 reserved for agentic orchestration. For this environment, a 60-minute time window was established for collecting episode summaries before triggering alpha value recommendations from the LLM Agent.
@@ 3.6.1 Baseline: SAC with adaptive $\alpha$
The baseline system employs the same RL-agent architecture described in Section 3.3, but uses the default adaptive $\alpha$ scheduling mechanism from Stable Baselines3 [14]. This implementation automatically adjusts the entropy coefficient during training according to the internal heuristics of the algorithm, without external intervention or episode-level performance feedback (Represented in Fig. 5).
@@ 4.2 On-Orbit Experiments (ISS)
ASTREA was subsequently deployed aboard the International Space Station (ISS) to validate its feasibility in a real operational environment (see Section 3.5.2). In the flight experiments, the maximum temperature was established at 57°C, a safe operational limit for the onboard equipment, with all experiments conducted during September 2025.
Two on-orbit experiments were conducted on this platform: a short-cycle configuration, and a longer one aligned with the orbital period.
4.2.1 Short-cycle experiment
For the first on-orbit configuration, the analysis window of the LLM agent was set to a conservative 15 minutes. This duration was selected as a prudential minimum considering the system’s measured inference latency: some queries required up to 8 minutes for completion (see Figure 7). Although a 10-minute window could have more closely matched the shortest stable orbital thermal cycles, the additional margin was intentionally introduced to ensure timing robustness and operational safety. The goal was to shorten the reasoning cycle as much as safely possible, considering the LLM’s inference latency and the fast thermal variations experienced in orbit.
In this configuration, the agentic system exhibited an increase in the number of thermal violations and also produced shorter episode durations than the RL-only baseline. CPU utilization decreased, reflecting inefficient resource allocation under these dynamic conditions (Table 2).
Table 2: Representative results from on-orbit short-cycle experiments aboard the ISS.
@@ 4.2.1 Short-cycle experiment
Avg. Episode Duration $\mathbf{449.77\pm 295.69}$ $363.76\pm 190.23$
Thermal Violations $\mathbf{66\pm 24.25}$ $82\pm 32.74$
Avg. CPU Usage (%) $\mathbf{40.17\pm 17.70}$ $22.86\pm 03.05$
The degraded performance profile can be partially explained by a temporal mismatch between the LLM decision cycle and the orbital thermal dynamics, combined with the stabilizing effect of the agentic policy on thermal control. Unlike terrestrial experiments, where ambient temperature evolves gradually and predictably, the ISS undergoes rapid and extreme thermal transitions during its 90-minute orbit—approximately 45 minutes of direct solar exposure followed by 45 minutes of eclipse. These sharp transitions often occurred while the LLM was still processing its previous cycle, rendering some $\alpha$ recommendations outdated at the moment of application.
As a result, certain updates introduced exploratory behavior during phases that required more conservative exploitation, while in other cases the agentic modulation successfully reduced the magnitude and frequency of thermal violations.
These observations motivated a second on-orbit configuration designed to align the reasoning cycle with the orbital period, aiming to test whether temporal synchronization could overcome the limitations identified in this setup.
4.2.2 Orbital-cycle experiment
For the second on-orbit configuration, the analysis window of the LLM agent was extended to 90 minutes to align with the ISS’s orbital period. This duration was chosen based on the observation that conditions in the immediate subsequent orbit are typically similar to those in the current one, allowing the system to leverage data from the previous window (orbit) as a proxy for estimating the optimal $\alpha$ value for the next one. By synchronizing the reasoning cycle with these predictable orbital dynamics the setup aimed to mitigate the temporal mismatches identified in the short-cycle experiment, ensuring that $\alpha$ recommendations remain relevant at the time of application while still accommodating the LLM’s inference latency.
In this configuration, the agentic system demonstrated substantial improvements across key metrics, achieving significantly longer episode durations and a marked reduction in thermal violations compared to the RL-only baseline. CPU utilization showed a 20.13% increase on average, reflecting enhanced resource efficiency without compromising thermal safety(Table 3).
Table 3: Representative results from on-orbit orbital-cycle experiments aboard the ISS.
Metric Baseline (Mean $\pm$ Std. Dev) Agentic (Mean $\pm$ Std. Dev)
Avg. Episode Duration $449.77\pm 295.69$ $\mathbf{1555.09\pm 823.57}$
Thermal Violations $66\pm 24.25$ $\mathbf{22.33\pm 9.07}$
Avg. CPU Usage (%) $40.17\pm 17.70$ $\mathbf{48.27\pm 11.49}$
The enhanced performance can be attributed to the improved alignment between the LLM’s decision cycle and the orbital thermal dynamics, which allowed for more timely and contextually appropriate $\alpha$ adjustments. This temporal synchronization enabled the agentic modulation to promote stable exploratory or exploitative behaviors that better anticipated the cyclical environmental shifts, reducing the risk of outdated recommendations. As a result, the system not only minimized thermal violations through proactive adaptation but also extended operational episodes by maintaining temperatures closer to the threshold without exceeding it. These findings validate the potential of orbit-aligned reasoning windows in overcoming latency constraints, providing a more robust framework for agentic supervision in dynamic LEO environments.
@@ 4.3 Comparative Discussion
The discrepancy between ground and on-orbit results underscores a critical limitation of agentic LLM-based systems in flight conditions: when the environment’s dominant timescale is comparable to or shorter than the LLM inference latency, semantic supervision can degrade performance.
In the ground experiments, ASTREA excelled, achieving a 67.12% increase in episode duration and a 58.48% reduction in thermal violations over the baseline during the first 4 hours, with stable CPU utilization, due to gradual thermal changes aligning with the LLM’s decision cycle.
However, the short-cycle on-orbit experiment (15-minute window) showed a 19.12% decrease in episode duration and a 24.24% increase in violations, as rapid thermal shifts rendered $\alpha$ recommendations outdated, often triggering inappropriate exploration during conservative phases.
The orbital-cycle experiment (90-minute window), aligning the reasoning cycle with the ISS orbit, reversed these trends, yielding a 245.75% increase in episode duration and a 66.17% reduction in violations, with a 20.13% increase in average CPU utilization. This success highlights that synchronizing the LLM’s analysis with environmental periodicity can mitigate latency issues, enabling effective supervision.
These findings affirm a key principle: LLMs are unsuitable for real-time intervention unless their latency is significantly shorter than the controlled process’s dynamics. The results provide practical design guidelines, emphasizing asynchronous, orbit-aligned architectures for future LEO autonomy systems.
@@ 5 Implications for Future Space Autonomy
The experimentation confirms that the integration of linguistic models for semantic supervision of other processes is feasible in flight-qualified equipment within the transformative New Space era. Agentic systems are thus positioned as a viable and, given their inherent characteristics, transformative approach to integrating control logic into traditional subsystems.
The deployment of ASTREA demonstrates how agentic systems can serve as supervisory layers for traditional control architectures. By combining linguistic reasoning with adaptive control, ASTREA exemplifies a scalable approach to onboard autonomy, capable of supporting fault detection, mission planning, and subsystem coordination.
@@ 6 Conclusions and Future Work
This work demonstrated the feasibility of deploying agentic systems on flight-qualified hardware to provide an upper layer of semantic supervision for autonomous edge control. By integrating a quantized Large Language Model (LLM) with a reinforcement learning (RL) agent, ASTREA showed that even compact linguistic models, despite their limited contextual capacity, can effectively complement adaptive controllers under strict hardware constraints. With appropriate optimization techniques, the hybrid architecture delivered measurable improvements in the thermal control agent’s performance, validating the potential of LLM-based reasoning to enhance long-term decision-making in spaceborne systems.
A single LLM agent was employed, exhibiting noticeable response-time limitations that would become even more critical in multi-agent configurations, where inference delays could generate bottlenecks, an aspect warranting further investigation. Likewise, the absence of dedicated accelerators on the current platform forced all inference to run on the CPU alongside other processes. The use of space-qualified accelerators could fundamentally transform performance, enabling larger models, shorter reasoning windows, and more frequent supervisory updates.
Although the LLM in this study focused on modulating the entropy coefficient ($\alpha$) of the RL model, future extensions could target other control parameters or incorporate adaptive reward shaping. Moreover, repeating the experiments with domain-specialized models in thermal management could strengthen contextual grounding and reduce dependence on prompt engineering. Overall, ASTREA establishes a concrete milestone toward embedding agentic reasoning in flight systems, demonstrating the technical readiness of semantic autonomy for next-generation space missions.